40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 1 17:07:06 2015, ericq, Update, LSC, IR beatnote confusion 
    Reply  Tue Sep 1 17:38:58 2015, ericq, Update, LSC, IR beatnote confusion 
       Reply  Wed Sep 2 13:44:04 2015, ericq, Update, LSC, IR beatnote confusion 
          Reply  Thu Sep 3 17:15:26 2015, ericq, Update, LSC, IR beatnote confusion GREENvIRbeats.pdf
             Reply  Fri Sep 4 09:36:48 2015, Steve, Update, LSC, IR beatnote confusion..... 
Message ID: 11556     Entry time: Tue Sep 1 17:07:06 2015     Reply to this: 11557
Author: ericq 
Type: Update 
Category: LSC 
Subject: IR beatnote confusion 

There has been some discussion here and there of using fiber coupled IR beats for ALS. A few weeks ago, and again today with Eric G, I poked around a bit with the fiber box Manasa set up for the FOL scheme. 

Somehow, the IR beatnote is ~1000 times smaller than expected, both with the Thorlabs fiber coupled PD and a fiber coupled NF 1611.

In essence, after the fiber combiner, there is on the order of hundreds of uW each of PSL and AUX X IR light. The output of the fiber from each source looks nice and gaussian. The DC output of the 1611 indicates that it is seeing the right level of light. The green beatnote exists with good SNR at twice the IR beat frequency, so we know that the IR beat isn't some junky modes beating. 

For the 1611, we would expect an RF signal of ~1mW*0.9A/W*700V/A -> .6V / +8dBm. Instead we see ~2mV / -40dBm.

Incidentally, there is some 20mV / -20dBm signal at ~400kHz, presumably from the green PDH modulation at ~200k. 

(The level out of the thorlabs PD is similarly tiny; it doesn't have a DC output though, so we don't know the DC power that the active surface really sees. Not that I expect it to be much different, but the NF just makes it easier to estimate.)

The only things that should be able to cause the beat to be smaller than expected from the power levels are mode matching and polarization matching. All the fibers are single mode, so mode matching should be effectively 100%. Maybe somthing fishy is happening with the polarizations, but they'd have to be really maliciously close to orthogonal to cause this level of mismatch. 

Maybe we just don't understand the splitter/combiners. Mysterious.

ELOG V3.1.3-